OpenStudy (anonymous):

Consider the line that passes through the point (-1, 6) and has a slope of -4. Part 1: Write the equation of this line using point-slope form. (2 points) Part 2: Using your equation from part 1, rewrite this equation in slope-intercept form. Make sure to show all of your work. (2 points) Part 3: Using your equation from part 2, rewrite this equation in standard form. Make sure to show all of your work.

OpenStudy (anonymous):

@AriPotta @Agent_Sniffles @radar HELP?

OpenStudy (anonymous):

@jlongSwag27 Any linear equations that you've known?

OpenStudy (anonymous):

No i just need help explaiining in detailing and everything because i dont understand this and this is for homework but i need to figure this out...

OpenStudy (anonymous):

Can you look up the POINT-SLOPE linear form?

OpenStudy (anonymous):

EXAMPLE @Chlorophyll y – y1 = m(x – x1) y – (–6) = (4)(x – (–1)) y + 6 = 4(x + 1) y + 6 = 4x + 4 y = 4x + 4 – 6 y = 4x – 2

OpenStudy (anonymous):

Exact the formula I want you to memorize! with ( x₁, y₁) is the given point @jlongSwag27 Now can you plug ( -1, 6) into the formula?